OpenStudy (anonymous):

A fair coin is flipped four times. What is the probability of getting heads at least once?

OpenStudy (anonymous):

didn't we just do this?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh you are right, 4 times. sorry

OpenStudy (anonymous):

we can list them or we can use a formula. your choice

OpenStudy (anonymous):

a formula

OpenStudy (anonymous):

at least once means 1 or 2 or 3 or 4, so not 0 the probability you get zero heads is \[(\frac{1}{2})^4\]so the probability you get at least one head is \[1-(\frac{1}{2})^4\]

OpenStudy (anonymous):

The sample space of the experiment is hhhh hhht hhtt httt hhth hthh htth htht thhh tthh ttth thth thht tttt thtt ttth so the number of times of getting heads at least once is 15 out of 16. But there is a better way. If you know that the probability of getting a heads at least once is mutually exclusive of getting all tails then P[heads at least once] = 1 - P[all tails] = 1 - 1/16 = 15/16

OpenStudy (anonymous):

of course zero heads mean all tails, i.e. you got t t t t

OpenStudy (anonymous):

satellite u r brainy